# Linear fractional relative risk aversion

### v1

@@ 3.2 Comparative statics results
Finally, assume that either $\alpha=0$ or $\beta=\frac{1}{\sigma}$ holds. In the former case, we have $p=\frac{m}{1-\beta}$, whereas in the latter case, we have $p=\frac{\sigma m}{\sigma-1}$. Hence, the markups are constant in $m$ and the prices are increasing in $m$ regardless of parameter values. Turning to quantities, we differentiate the first-order condition $u^{\prime}(q)=\nu p$ with respect to $m$ to obtain $u^{\prime\prime}(q)\frac{\partial q}{\partial m}=\nu\frac{\partial p}{\partial m}$. Since $u^{\prime\prime}(q)<0$ and $\nu>0$, $\frac{\partial q}{\partial m}$ and $\frac{\partial p}{\partial m}$ must have opposite signs, thus implying that the quantities are decreasing in $m$ regardless of parameter values. This completes the proof. $\square$
Table 7: Comparative statics.
Case 1: $\alpha>0$.
markups prices quantities
Case 2: $\alpha<0$.
markups prices quantities
Notes: The case with $\alpha=0$ corresponds to the CES and requires that $\beta\in(0,1)$. The case with $1-\beta\sigma=0$ also corresponds to the CES and requires that $\sigma>1$. If $\alpha=0$ or $\beta=\frac{1}{\sigma}$, then the price is given by $p(m)=\frac{m}{1-\beta}$ or $p(m)=\frac{\sigma m}{\sigma-1}$, respectively, and thus $\frac{\partial}{\partial m}\bigl{(}\frac{1}{\mu}\bigr{)}=0$, $\frac{\partial}{\partial m}\bigl{(}\frac{m}{\mu}\bigr{)}>0$, and $\frac{\partial}{\partial m}\bigl{(}q(m)\bigr{)}=\frac{\partial}{\partial m}\bigl{(}(u^{\prime})^{-1}(\nu p(m))\bigr{)}<0$. Since these two cases are well known, we omit them from the table.
It is worth emphasizing that our parametric approach allows us to take the model to data and estimate the three parameters $\{\alpha,\beta,\sigma\}$ of the LFRRA utility function. These estimates can then be used to test whether the RRA is increasing, decreasing, and constant, and thus whether markups are decreasing, increasing, and constant with respect to marginal costs. We present these results in the next section.
Fitting the monopolistic competition model to data
We quantify the model using the LFRRA specification (1) and the first-order condition for profit maximization (21), which implies $1-\mu(\omega)=\frac{\alpha q(\omega)+\beta}{\alpha\sigma q(\omega)+1}$. Using observed firm-level markups $\frac{1}{\mu(\omega)}$ and quantities $q(\omega)$, we estimate the values of $\alpha$, $\beta$, and $\sigma$ by minimizing the following residual sum of squares (RSS):
subject to the theoretical constraints $\alpha q(\omega)+\beta>0$, $\alpha\sigma q(\omega)+1>0$, and $0<\frac{\alpha q(\omega)+\beta}{\alpha\sigma q(\omega)+1}\leq 1$ (see Appendix A for more details). We solve the minimization problem (38) at the sectoral and aggregate levels to obtain the fitted values $\{\widehat{\alpha},\widehat{\beta},\widehat{\sigma}\}$ that are consistent with our LFRRA preferences. We also verify that the second-order conditions (24) evaluated at $\{\widehat{\alpha},\widehat{\beta},\widehat{\sigma}\}$ are satisfied for all quantities $q(\omega)\geq 0$.
To conduct statistical inference, we construct confidence intervals for $\{\alpha,\beta,\sigma\}$ by bootstrapping as follows. We first randomly draw, with replacement, a sample of markup and quantity observations from the data such that the generated and observed sample sizes are identical. We then compute $\{\widetilde{\alpha},\widetilde{\beta},\widetilde{\sigma}\}$ by solving the minimization problem (38) using the procedure explained in Appendix A. We repeat this process 200 times at the sectoral and aggregate levels to obtain the distributions of parameters $\{\widetilde{\alpha}_{b},\widetilde{\beta}_{b},\widetilde{\sigma}_{b}\}_{b=1}^{200}$, which allow us to construct the $95\%$ confidence interval using the 2.5th and 97.5th percentiles of the distribution of each parameter, respectively.
We further consider two alternative cases in Appendix A. First, we use the specification in Corollary 2, which encompasses the alternative CREMR specification by Mrázová et al. (2021). Second, we consider non-linear fractional RRA as in Klenow and Willis (2016) and Edmond et al. (2023) with $\frac{qu^{\prime\prime}(q)}{u^{\prime}(q)}=\frac{1}{\overline{\sigma}q^{-\varepsilon/\overline{\sigma}}}$ (in the latter paper’s notation). We find that the value of the RRS for each sector and the aggregate economy in Table 8 is smaller than the corresponding values in Tables A.3 and A.4, thus showing that those specifications fit the data from De Loecker et al. (2016) less well than our LFRRA.
Table 8: Sectoral and aggregate estimates of the LFRRA parameters.
[…]
17 Textiles, apparel 22.692 0.127 0.257 5.760 0.084 n.a. CRRA 1534
24 Chemicals 35.152 118.472 0.240 2.774 -0.120 39.523 IRRA 1684
[…]
27 Basic metals 20.660 17.830 0.281 5.304 0.093 -8.756 DRRA 988
[…]
31 Electrical machinery 8.182 0.426 0.280 2.219 -0.170 n.a. CRRA 495
[…]
Notes: We use markup and quantity data from De Loecker et al. (2016). We restrict ourselves to single-product firms and products for which the markup exceeds one. Following De Loecker et al. (2016), we trim the top- and bottom-3% of quantities and markups by sector. We run 200 bootstrap replications with replacement, and the 95% confidence intervals are provided below the coefficients. RSS denotes the residual sum of squares, i.e., the value of our objective function in (38). Obs stands for the number of observations. IRRA, DRRA, and CRRA stand for increasing, decreasing, and constant relative risk aversion, respectively. Details on the numerical procedure are provided in Appendix A. All values in excess of $10^{4}$ are reported using scientific notation. means that the derivative of RRA is zero when $\widehat{\alpha}$ or $\widehat{\beta}-{1}/{\widehat{\sigma}}$ is not different from zero and is therefore not applicable.
Implicitly additive preferences
Consider the implicitly additively separable utility function $U$, defined as $\int_{\Omega}\Upsilon\bigl{(}\frac{q(\omega)}{U}\bigr{)}{\rm d}\omega=1$, where $\Omega$ denotes the set of varieties, $q(\omega)\geq 0$ is consumption of variety $\omega$, and $\Upsilon$ is an increasing and concave sub-function (Kimball, 1995; Mrázová et al., 2021). Minimizing expenditure $\int_{\Omega}p(\omega)q(\omega){\rm d}\omega$ subject to the utility constraint $\int_{\Omega}\Upsilon\bigl{(}\frac{q(\omega)}{U}\bigr{)}{\rm d}\omega=1$ gives the first-order condition $p(\omega)=\frac{\nu^{\rm imp}}{U}\Upsilon^{\prime}\bigl{(}\frac{q(\omega)}{U}\bigr{)}$, where $\nu^{\rm imp}>0$ is the Lagrange multiplier.
Using this condition, the profit maximization problem in the monopolistic competition model in Section 3 can be replaced with $\max_{q}\bigl{[}\frac{\nu^{\rm imp}}{U}\Upsilon^{\prime}\bigl{(}\frac{q}{U}\bigr{)}-m\bigr{]}LU\frac{q}{U}-f$. Let $z=\frac{q}{U}$ denote scaled quantities. We can then rewrite the problem as $\max_{z}\bigl{[}\frac{\nu^{\rm imp}}{U}\Upsilon^{\prime}(z)-m\bigr{]}LUz-f$. Since there is a continuum of firms, no firm can affect $\nu^{\rm imp}$ and $U$. Thus, the first-order condition for profit maximization is given by $\frac{\nu^{\rm imp}}{U}\Upsilon^{\prime\prime}(z)LUz+\bigl{[}\frac{\nu^{\rm imp}}{U}\Upsilon^{\prime}(z)-m\bigr{]}LU=0$. The profit maximizing price $p(m)$ hence satisfies:
$\frac{p(m)-m}{p(m)}=1-\mu(m)=-\frac{z(m)\Upsilon^{\prime\prime}(z(m))}{\Upsilon^{\prime}(z(m))},$ (39)
where $\mu(m)$ denotes the inverse of the markup. The second-order condition (SOC) for profit maximization is given by
$\textstyle\frac{\nu^{\rm imp}}{U}\Upsilon^{\prime\prime\prime}(z)LUz+\frac{\nu^{\rm imp}}{U}\Upsilon^{\prime\prime}(z)LU+\frac{\nu^{\rm imp}}{U}\Upsilon^{\prime\prime}(z)LU$ $=$ $\nu^{\rm imp}\Upsilon^{\prime\prime\prime}(z)Lz+2\nu^{\rm imp}\Upsilon^{\prime\prime}(z)L$
$=$ $\textstyle-L\nu^{\rm imp}\Upsilon^{\prime\prime}(z)\left[-\frac{z\Upsilon^{\prime\prime\prime}(z)}{\Upsilon^{\prime\prime}(z)}-2\right]<0.$
Since $L\nu^{\rm imp}\Upsilon^{\prime\prime}(z)>0$ , the SOC reduces to $\frac{z\Upsilon^{\prime\prime\prime}(z)}{\Upsilon^{\prime\prime}(z)}<2$.
As in Section 2, we assume that $\frac{z(m)\Upsilon^{\prime\prime}(z(m))}{\Upsilon^{\prime}(z(m))}$ is of the linear fractional form, $\frac{\alpha z(m)+\beta}{\alpha\sigma z(m)+1}$, so that (39) becomes
$1-\mu(m)=\frac{\alpha z(m)+\beta}{\alpha\sigma z(m)+1}.$ (40)
We can fit (40) to data in the same way as in Section 4 by noting that $1-\mu(m)=\frac{\alpha z(m)+\beta}{\alpha\sigma z(m)+1}=\frac{\alpha\frac{q(m)}{U}+\beta}{\alpha\sigma\frac{q(m)}{U}+1}=\frac{\widetilde{\alpha}q(m)+\beta}{\widetilde{\alpha}\sigma q(m)+1}$, where $\widetilde{\alpha}=\frac{\alpha}{U}$, $\beta$, and $\sigma$ are the parameters to be estimated.
Conclusion
In this paper we first characterize the LFRRA family of utility functions in terms of the Gauss hypergeometric functions. This family, which is parametrized by $\{\alpha,\beta,\sigma\}$, nests the CES, CARA, quadratic, log, HARA, CREMR, incomplete gamma, incomplete beta, and exponential integral utility functions as special cases. We then embed this family into a monopolistic competition model and derive the price equilibrium. We show that the inverse markup function is expressed as the generalized Lambert $W$ function, ${\cal W}(x,\alpha,\beta,\sigma)$, which reduces to the original Lambert $W$ function when $\beta$ and $\sigma$ are set to zero. We finally take the relationship between the inverse markup and the LFRRA to firm-level data and estimate $\{\widehat{\alpha},\widehat{\beta},\widehat{\sigma}\}$. The estimates are used to test whether each sector displays IRRA, DRRA, or CRRA, which in turn determines whether markups are decreasing, increasing, or constant in marginal costs. Our results show that all cases are present in the data, thus highlighting the importance of taking a flexible parametric approach.
We believe that the LFRRA will be useful to fill the gap between the empirical literature on markups, which shows that markups are heterogeneous across firms in each sector, and quantitative trade and urban models, where markups are typically assumed to be constant. Extending the LFRRA family to such general equilibrium analysis will require to replace the inverse markup function or the Lambert $W$ function in Behrens et al. (2014, 2017, 2020) with our generalized Lambert $W$ function. Observe that the benefit of using the LFRRA is not limited to the price equilibrium in trade and urban models. It would also spur the development of qualitative and quantitative analysis in various other contexts such as economic growth and the economics of uncertainty, where the RRA matters.
Appendix Appendix A Details on the empirical implementation
@@ Conclusion
We solve the minimization problem
$\min_{\{\alpha,\beta,\sigma\}}\ \text{RSS}=\sum_{\omega=1}^{\Omega}\left[1-\mu(\omega)-\frac{\alpha q(\omega)+\beta}{\alpha\sigma q(\omega)+1}\right]^{2}$ (A.1)
subject to the theoretical constraints $\alpha q(\omega)+\beta>0$, $\alpha\sigma q(\omega)+1>0$, and $0<\frac{\alpha q(\omega)+\beta}{\alpha\sigma q(\omega)+1}\leq 1$. As shown by (A.1), conditional on observed markups $\mu(\omega)$ and quantities $q(\omega)$, only the product $\alpha q(\omega)$ matters for the LFRRA. Without loss of generality, we therefore scale quantities by the mean quantity $\overline{q}$ in the sample and use $q(\omega)/\overline{q}$ as our measure of quantities. Doing so improves the numerical computations since quantities are extremely skewed. In what follows we refrain from interpreting the magnitude of $\alpha$ as it depends on the scaling.
Following Arkolakis et al. (2019), we split the estimation procedure into an inner loop and an outer loop. In the inner loop, we fix the value of $\alpha$ and minimize the RSS in (A.1) with respect to $\beta$ and $\sigma$, which yields $\widehat{\beta}(\alpha)$ and $\widehat{\sigma}(\alpha)$. In the outer loop, we do a grid search for the value $\widehat{\alpha}$ that minimizes the RSS in (A.1) conditional on $\widehat{\beta}(\alpha)$ and $\widehat{\sigma}(\alpha)$.
@@
For the sake of completeness, we also re-estimated the values of $\alpha$ and $\sigma$ by constraining $\beta$ a priori to be either 0 or $1$. This allows us to check that our procedure has ‘correct’ numerical behavior at the corners. Reassuringly, as should be the case, the RSS is never smaller than the one we estimate in the unconstrained model. We show these results in Tables A.1 and A.2.
Next, although by Assumption 2 the case with $\alpha\sigma<0$ is not nested in the CREMR specification (13) in Theorem 2, we can estimate this case using the specification (20) in Corollary 2, which requires the restrictions, $\alpha=-\frac{1}{\gamma\sigma}$, $\sigma>0$, $q_{\min}>\gamma$, and $\gamma>0$. The results are given in Table A.3. We do not find any case in which the specification (20) in Corollary 2 yields a smaller RSS than the specification (13) in Theorem 2 in Table 8.363636Our exercise differs from that in Mrázová et al. (2021) who estimate preference and productivity distribution parameters for a given pair of utility and productivity distribution functions. In contrast, we estimate three preference parameters in the LFRRA without specifying the productivity distribution function. t
Finally, we consider non-linear fractional RRA as in Klenow and Willis (2016) and Edmond et al. (2023) with $\frac{qu^{\prime\prime}(q)}{u^{\prime}(q)}=\frac{1}{\overline{\sigma}q^{-\varepsilon/\overline{\sigma}}}$ (in the latter paper’s notation). Comparing Tables 8 and A.4 shows that the RRS for each sector and the aggregate economy is larger for the non-linear fractional RRA than for our LFRRA, meaning that the former specification fits the data from De Loecker et al. (2016) less well than the latter.
@@
Table A.1: Sectoral and aggregate estimates (LFRRA constrained by $\beta=0$).
@@
Table A.2: Sectoral and aggregate estimates (LFRRA constrained by $\beta=1$).
@@
Table A.3: Sectoral and aggregate estimates (Mrázová et al., 2021).
@@
Table A.4: Sectoral and aggregate estimates (Klenow and Willis, 2016; and Edmond et al., 2023).
@@
Sector Name RSS $\widehat{\varepsilon}$ $\widehat{\overline{\sigma}}$ Type Obs 15 Food products and beverages 54.487 -0.012 2.857 CRRA 1044 [0.041,0.014][-0.041,0.014] $[2.663,3.076]$ 17 Textiles, apparel 22.833 -0.015 4.021 CRRA 1534 [0.073,0.026][-0.073,0.026] $[3.730,4.335]$ 21 Paper and paper products 6.541 0.049 2.988 CRRA 546 [0.016,0.105][-0.016,0.105] $[2.725,3.297]$ 24 Chemicals 35.971 0.114 2.636 IRRA 1684 $[0.091,0.133]$ $[2.494,2.813]$ 25 Rubber and plastic 11.898 0.004 3.258 CRRA 676 [0.031,0.033][-0.031,0.033] $[3.026,3.511]$ 26 Nonmetallic mineral products 24.169 0.042 1.657 IRRA 744 $[0.025,0.057]$ $[1.576,1.741]$ 27 Basic metals 20.753 -0.291 5.524 DRRA 988 [0.427,0.162][-0.427,-0.162] $[5.032,6.056]$ 28 Fabricated metal products 7.150 -0.051 3.858 DRRA 445 [0.120,0.002][-0.120,-0.002] $[3.345,4.401]$ 29 Machinery and equipment 15.351 -0.033 3.585 DRRA 745 [0.085,0.004][-0.085,-0.004] $[3.334,3.952]$ 31 Electrical machinery 8.443 0.012 3.300 CRRA 495    and communications [0.009,0.031][-0.009,0.031] $[3.057,3.580]$ 34 Motor vehicles, trailers 8.625 -0.080 3.434 DRRA 331 [0.133,0.026][-0.133,-0.026] $[3.100,3.796]$ 0 All sectors pooled 274.005 0.037 2.669 IRRA 9232 $[0.030,0.043]$ $[2.579,2.795]$ Notes: We use markup and quantity data from De Loecker et al. (2016). We restrict ourselves to single-product firms and products for which the markup exceeds one. Following De Loecker et al. (2016), we trim the top- and bottom-3% of quantities and markups by sector. We run 200 bootstrap replications with replacement, and the 95% confidence intervals are provided below the coefficients. RSS denotes the residual sum of squares, i.e., the value of our objective function in (38). Obs stands for the number of observations. Details on the numerical procedure are provided in Appendix A. All values in excess of $10^{4}$ are reported using scientific notation.

### v2

@@ Appendix B.3 Proof of Proposition 5
Finally, assume that either $\alpha=0$ or $\beta=\frac{1}{\sigma}$ holds. In the former case, we have $p=\frac{m}{1-\beta}$, whereas in the latter case, we have $p=\frac{\sigma m}{\sigma-1}$. Hence, the markups are constant in $m$ and the prices are increasing in $m$ regardless of parameter values. Turning to quantities, we differentiate the first-order condition $u^{\prime}(q)={\lambda}p$ with respect to $m$ to obtain $u^{\prime\prime}(q)\frac{\partial q}{\partial m}={\lambda}\frac{\partial p}{\partial m}$. Since $u^{\prime\prime}(q)<0$ and ${\lambda}>0$, $\frac{\partial q}{\partial m}$ and $\frac{\partial p}{\partial m}$ must have opposite signs, thus implying that the quantities are decreasing in $m$ regardless of parameter values. This completes the proof. $\square$
Numerical implementation
@@
We solve the minimization problem (36) subject to the theoretical constraints $\alpha q(\omega)+\beta>0$, $\alpha\sigma q(\omega)+1>0$, and $0<\frac{\alpha q(\omega)+\beta}{\alpha\sigma q(\omega)+1}\leq 1$. As shown by (36), conditional on observed markups $\mu(\omega)$ and quantities $q(\omega)$, only the product $\alpha q(\omega)$ matters for the LFRRA. Without loss of generality, we therefore scale quantities by the mean quantity $\overline{q}$ in the sample and use $q(\omega)/\overline{q}$ as our measure of quantities. Doing so improves the numerical computations since quantities are extremely skewed. In what follows we refrain from interpreting the magnitude of $\alpha$ as it depends on the scaling.
Following Arkolakis et al. (2019), we split the estimation procedure into an inner loop and an outer loop. In the inner loop, we fix the value of $\alpha$ and minimize the RSS in (36) with respect to $\beta$ and $\sigma$, which yields $\widehat{\beta}(\alpha)$ and $\widehat{\sigma}(\alpha)$. In the outer loop, we do a grid search for the value $\widehat{\alpha}$ that minimizes the RSS in (36) conditional on $\widehat{\beta}(\alpha)$ and $\widehat{\sigma}(\alpha)$.
@@
For the sake of completeness, we also re-estimated the values of $\alpha$ and $\sigma$ by constraining $\beta$ a priori to be either 0 or $1$. This allows us to check that our procedure has ‘correct’ numerical behavior at the corners. Reassuringly, as should be the case, the RSS is never smaller than the one we estimate in the unconstrained model. We show these results in Tables C.1 and C.2.
Next, although by Assumption 2 the case with $\alpha\sigma<0$ is not nested in the CREMR specification (10) in Theorem 2, we can estimate this case using the specification (15) in Corollary 2, which requires the restrictions, $\alpha=-\frac{1}{\gamma\sigma}$, $\sigma>0$, $q_{\min}>\gamma$, and $\gamma>0$. The results are given in Table C.3. We do not find any case in which the specification (15) in Corollary 2 yields a smaller RSS than the specification (10) in Theorem 2 in Table 7.434343Our exercise differs from that in Mrázová et al. (2021) who estimate preference and productivity distribution parameters for a given pair of utility and productivity distribution functions. In contrast, we estimate three preference parameters in the LFRRA without specifying the productivity distribution function.
Finally, we consider two non-linear fractional forms introduced in Section 4.2. One is $\frac{qu^{\prime\prime}(q)}{u^{\prime}(q)}=\frac{\gamma}{q^{\frac{k-1}{k}}+\gamma}$ as in Mrázová and Neary (2017) and the other is $\frac{\theta\Upsilon^{\prime\prime}(\theta)}{\Upsilon^{\prime}(\theta)}=\frac{1}{\overline{\sigma}\theta^{-\frac{\varepsilon}{\overline{\sigma}}}}$ with $\theta=\frac{q}{U}$ as in Klenow and Willis (2016) and Edmond et al. (2023). Comparing Tables C.4 and C.5 and Table 7 shows that the RRS for each sector and the aggregate economy is larger for these two non-linear fractional forms than for our LFRRA, meaning that the former specifications fit the data from De Loecker et al. (2016) less well than the latter.
@@
Table C.1: Sectoral and aggregate estimates (LFRRA constrained by $\beta=0$).
@@
Table C.2: Sectoral and aggregate estimates (LFRRA constrained by $\beta=1$).
@@
Table C.3: Sectoral and aggregate estimates (CREMR in Mrázová et al., 2021).
@@
Table C.4: Sectoral and aggregate estimates (CPPT in Mrázová and Neary, 2017).
@@
Sector Name RSS $\widehat{\gamma}$ $\widehat{k}$ Type Obs 15 Food products and beverages 54.487 0.538 1.007 CRRA 1044 $[0.480;0.601]$ $[0.991;1.021]$ 17 Textiles, apparel 22.833 0.331 1.005 CRRA 1534 $[0.287;0.366]$ $[0.986;1.023]$ 21 Paper and paper products 6.540 0.504 0.976 CRRA 546 $[0.006;0.576]$ $[0.276;1.007]$ 24 Chemicals 35.911 0.613 0.939 IRRA 1684 $[0.552;0.675]$ $[0.924;0.954]$ 25 Rubber and plastic 11.898 0.443 0.998 CRRA 676 $[0.395;0.494]$ $[0.984;1.014]$ 26 Nonmetallic mineral products 24.242 1.467 0.952 IRRA 744 $[1.320;1.633]$ $[0.935;0.973]$ 27 Basic metals 20.759 0.219 1.074 CRRA 988 $1.597\times 10^{-8};0.247] [$0.117;1.103][0.117;1.103] 28 Fabricated metal products 7.155 0.353 1.018 DRRA 445 $[0.292;0.426]$ $[1.000;1.038]$ 29 Machinery and equipment 15.351 0.387 1.013 DRRA 745 $[0.337;0.429]$ $[1.002;1.031]$ 31 Electrical machinery 8.444 0.433 0.995 CRRA 495    and communications $[0.389;0.482]$ $[0.986;1.003]$ 34 Motor vehicles, trailers 8.611 0.404 1.039 IRRA 331 $[0.343;0.472]$ $[1.013;1.061]$ 0 All sectors pooled 273.993 0.596 0.980 IRRA 9232 $7.326\times 10^{-11};0.634] [$0.122;0.985][0.122;0.985] Notes: We use markup and quantity data from De Loecker et al. (2016). We restrict ourselves to single-product firms and products for which the markup exceeds one. Following De Loecker et al. (2016), we trim the top- and bottom-3% of quantities and markups by sector. We run 200 bootstrap replications with replacement, and the 95% confidence intervals are provided below the coefficients. RSS denotes the residual sum of squares, i.e., the value of our objective function in (36). Obs stands for the number of observations. Details on the numerical procedure are provided in Appendix C. All values in excess of $10^{4}$ are reported using scientific notation.
Table C.5: Sectoral and aggregate estimates (Klenow and Willis, 2016; and Edmond et al., 2023).
Sector Name RSS $\widehat{\varepsilon}$ $\widehat{\overline{\sigma}}$ ${\theta\Upsilon^{\prime\prime}(\theta)}/{\Upsilon^{\prime}(\theta)}$ Obs 15 Food products and beverages 54.487 -0.012 2.857    constant 1044 [0.041,0.014][-0.041,0.014] $[2.663,3.076]$ 17 Textiles, apparel 22.833 -0.015 4.021    constant 1534 [0.073,0.026][-0.073,0.026] $[3.730,4.335]$ 21 Paper and paper products 6.541 0.049 2.988    constant 546 [0.016,0.105][-0.016,0.105] $[2.725,3.297]$ 24 Chemicals 35.971 0.114 2.636    increasing 1684 $[0.091,0.133]$ $[2.494,2.813]$ 25 Rubber and plastic 11.898 0.004 3.258    constant 676 [0.031,0.033][-0.031,0.033] $[3.026,3.511]$ 26 Nonmetallic mineral products 24.169 0.042 1.657    increasing 744 $[0.025,0.057]$ $[1.576,1.741]$ 27 Basic metals 20.753 -0.291 5.524    decreasing 988 [0.427,0.162][-0.427,-0.162] $[5.032,6.056]$ 28 Fabricated metal products 7.150 -0.051 3.858    decreasing 445 [0.120,0.002][-0.120,-0.002] $[3.345,4.401]$ 29 Machinery and equipment 15.351 -0.033 3.585    decreasing 745 [0.085,0.004][-0.085,-0.004] $[3.334,3.952]$ 31 Electrical machinery 8.443 0.012 3.300    constant 495    and communications [0.009,0.031][-0.009,0.031] $[3.057,3.580]$ 34 Motor vehicles, trailers 8.625 -0.080 3.434    decreasing 331 [0.133,0.026][-0.133,-0.026] $[3.100,3.796]$ 0 All sectors pooled 274.005 0.037 2.669    increasing 9232 $[0.030,0.043]$ $[2.579,2.795]$ Notes: We use markup and quantity data from De Loecker et al. (2016). We restrict ourselves to single-product firms and products for which the markup exceeds one. Following De Loecker et al. (2016), we trim the top- and bottom-3% of quantities and markups by sector. We run 200 bootstrap replications with replacement, and the 95% confidence intervals are provided below the coefficients. RSS denotes the residual sum of squares, i.e., the value of our objective function in (36). Obs stands for the number of observations. Details on the numerical procedure are provided in Appendix C. All values in excess of $10^{4}$ are reported using scientific notation.